# On decomposability and subdifferential of the tensor nuclear norm

### v1

@@ 6.2 Unique optimality
$\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\bigl\|\operatorname{p}_{\mathbb{L}}(\mathbf{D})-\mathbf{Z}\bigr\|_{2}\leq\frac{\lambda}{8},~\bigl\|\operatorname{p}_{\mathbb{L}^{\perp}}(\mathbf{D})\bigr\|_{\sigma}<\frac{1}{2},~\bigl\|\operatorname{p}_{\mathbb{I}(\mathbf{S})}(\mathbf{D})-\lambda\mathbf{E}\bigr\|_{2}\leq\frac{\lambda}{8},\text{ and }\bigl\|\operatorname{p}_{\mathbb{I}^{\perp}(\mathbf{S})}(\mathbf{D})\bigr\|_{\infty}<\frac{\lambda}{2}.$
@@
The most important result in this part is to bound the tail probability of $\bigl\|\operatorname{p}_{\mathbb{L}}(\operatorname{p}_{\mathbb{I}^{d}}-q^{-1}\operatorname{p}_{\mathbb{I}})\operatorname{p}_{\mathbb{L}}\bigr\|$, i.e., Proposition 6.6. To achieve this, we treat a projection onto a subspace in $\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$, which is itself a linear operator, as a $\prod_{k=1}^{d}n_{k}$ by $\prod_{k=1}^{d}n_{k}$ matrix, and then apply the matrix Bernstein inequality [55, Theorem 1.4]: If $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{m}\in\mathbb{R}^{n\times n}$ are independent and self-adjoint random matrices such that $\operatorname{Ex}\mathbf{X}_{i}=\mathbf{O}$ and $\text{emax}(\mathbf{X}_{i})\leq s$ almost surely for any $i\in[m]$, then for any $t\geq 0$,
@@ 6.4 Dual certificate: Low-rank part via the golfing scheme
In this subsection, we construct the low-rank part of the dual certificate via the golfing scheme [22, 11]; let us start with a couple of probability bounds before introducing the scheme. The first one is a direct consequence of [64, Theorem 2.1]: If $d$th order tensors $\mathbf{X}\in\mathbb{R}^{n\times n\times\dots\times n}$ and $\mathbf{Y}\in\{0,1\}^{n\times n\times\dots\times n}$ with $\mathbf{X}$ given and $\mathbb{I}(\mathbf{Y})\sim\operatorname{Bernoulli}(q)$ for $q\geq\theta_{1}\frac{\ln n}{n}$, then
@@
By applying the scalar Bernstein inequality [56, Theorem 2.8.4], we have
@@
As result, for a sufficiently large $\kappa_{3}>0$,
@@
holds with high probability, where the last equality holds because $\operatorname{p}_{\mathbb{L}}(\mathbf{Z}_{j-1})-\mathbf{Z}\in\mathbb{L}$ and the second inequality is to due Corollary 6.9. The proof is then completed by an overall union bound. $\Box$
@@ 6.5 Dual certificate: Sparse part via the least squares method
In order to show (32), the existence of relevant $\epsilon$-nets is required. In particular, it is stated in [56, Corollary 4.2.13] that for any $\epsilon>0$, there exists an $\epsilon$-net of $\mathbb{S}^{n}$ whose cardinality is no more than $(1+\frac{2}{\epsilon})^{n}$. $\Box$
@@
This, together with [56, Proposition 2.5.2], further implies that
@@
The result then follows directly by combining the above with [54, Lemma 1] and [54, Theorem 1]. $\Box$
@@
where $\mathbb{E}_{k}$’s are the $\epsilon$-nets used in (32). For any given $\mathbb{I}(\mathbf{S})$, the nonzero entries of $\mathbf{E}$ are i.i.d. symmetric Bernoulli random variables (taking $\pm 1$ with equal probability), according to Assumption 6.1. Hence, by Hoeffding’s inequality [56, Theorem 2.2.2], we have
@@
For any given $\mathbb{I}(\mathbf{S})$, the nonzero entries of $\mathbf{E}$ are i.i.d. symmetric Bernoulli random variables. By the two-sided version of Hoeffding’s inequality [56, Theorem 2.2.5],
@@ 6.6 Putting everything together
We arrive at the following result by combining the results of previous two subsections with an overall union bound. This, together with Corollary 6.8 and Lemma 6.4, directly shows Theorem 6.2.
@@ 7 Concluding remarks
In this paper, we systematically study the decomposability and the subdifferential of the tensor nuclear norm. We show that the tensor nuclear norm is decomposable over a pair of subspaces that have at least two disjoint modes, naturally generalizing the result for the matrix case. The same property holds for the tensor spectral norm as well. Based on the decomposability, we propose novel inclusions of the subdifferential of the tensor nuclear norm. They strictly enlarge the inclusion proposed in [62, Theorem 1], the only known subdifferential inclusion for tensors of an arbitrary order. While various bounds for the subdifferential of the tensor nuclear norm and several interesting examples suggest that there is no general way to explicitly characterize the subdifferential of the tensor nuclear norm in general as in the matrix case, we examine subgradients of the tensor nuclear norm in all relevant subspaces and estimate their inner and outer approximations.
@@
In fact, the conjecture holds true if the factor $\ln^{d-2}{n_{d}}$ in Corollary 6.9 can be removed. This has already been conjectured similarly in [64, Section 11]. If Conjecture 7.1 is true, then the conditions for exact recovery of the tensor robust PCA of every order, including the matrix case, would become identical.
@@
Apart from practical applications, our developments may facilitate the study of a variety of other more in-depth properties and implications of the tensor nuclear norm. For example, they may help to deduce the $C^{2}$-cone reducibility [6, Definition 3.135], subdifferential metric subregularity [18, Section 3.H], and twice epi-differentiability [53, Definition 13.6(b)] of the tensor nuclear norm; see, e.g., [16, Proposition 3.2][63, Proposition 11] and [16, Proposition 3.8], and [24, Corollary 3.6] and [43], respectively for their matrix counterparts. Besides, the developments can also be useful to understand the neural collapse [48] in training tensor-based neural networks and the relations between various nuclear-norm-regularized tensor optimization problems and their Burer-Monteiro reparameterizations [9, 10]; see, e.g., [65, Appendix C] and [34, 8, 42, 47], respectively for their matrix counterparts.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China [Grants 72394364, 72394360, and 72171141]. The first author would like to express his sincere gratitude to Professor Simai He for his generous subsidy support when the first author was studying at School of Information Management and Engineering, Shanghai University of Finance and Economics, where part of the work by the first author was conducted.
@@ Appendix A Computing tensor spectral norms
Since $\mathbf{Z}+\mathbf{X}(t)$ is a subtensor of $\mathbf{U}(t)$, it is easy to see (or by [35, Theorem 3.1]) that
@@
Proof. Since $\mathbf{X}(t)$ is a symmetric tensor, it follows from Banach’s theorem (see e.g., [2] and [14, Corollary 4.2]) that
@@
We remark that $\bigl\|\mathbf{X}(\frac{1}{\sqrt{3}})\bigr\|_{\sigma}$ has been calculated in [20, Lemma 6.2] and can be easily extended to $\bigl\|\mathbf{X}(t)\bigr\|_{\sigma}$ for any $t$. $\bigl\|\mathbf{Z}+\mathbf{X}(-1)\bigr\|_{\sigma}$ has also been calculated in [20, Lemma 6.1]. In fact, the tensor $\mathbf{Z}+\mathbf{X}(-1)$ is known as an orthogonal tensor in the literature; see [36, Theorem 3.5] for more details. The result of $\bigl\|\mathbf{Z}+\mathbf{X}(t)\bigr\|_{\sigma}$ in Lemma A.3 provides a better understanding of the generalization of such a tensor.
@@ Appendix C Subdifferential of the nuclear norm for fourth-order tensors
The following result generalizes [61, Lemma 1] from $d=3$ to $d=4$.

### v2

@@ 6.2 Unique optimality
$\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\bigl\|\operatorname{p}_{\mathbb{L}}(\mathbf{D})-\mathbf{Z}\bigr\|_{2}\leq\frac{\lambda}{8},~\bigl\|\operatorname{p}_{\mathbb{L}^{\perp}}(\mathbf{D})\bigr\|_{\sigma}<\frac{1}{2},~\bigl\|\operatorname{p}_{\mathbb{I}(\mathbf{S})}(\mathbf{D})-\lambda\mathbf{E}\bigr\|_{2}\leq\frac{\lambda}{8},\text{ and }\bigl\|\operatorname{p}_{\mathbb{I}^{\perp}(\mathbf{S})}(\mathbf{D})\bigr\|_{\infty}<\frac{\lambda}{2}.$ (27)
@@
The most important result in this part is to bound the tail probability of $\bigl\|\operatorname{p}_{\mathbb{L}}(\operatorname{p}_{\mathbb{I}^{d}}-q^{-1}\operatorname{p}_{\mathbb{I}})\operatorname{p}_{\mathbb{L}}\bigr\|$, i.e., Proposition 6.6. To achieve this, we treat a projection onto a subspace in $\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$, which is itself a linear operator, as a $\prod_{k=1}^{d}n_{k}$ by $\prod_{k=1}^{d}n_{k}$ matrix, and then apply the matrix Bernstein inequality [63, Theorem 1.4]: If $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{m}\in\mathbb{R}^{n\times n}$ are independent and self-adjoint random matrices such that $\operatorname{Ex}\mathbf{X}_{i}=\mathbf{O}$ and $\text{emax}(\mathbf{X}_{i})\leq s$ almost surely for any $i\in[m]$, then for any $t\geq 0$,
@@ 6.4 Dual certificate: Low-rank part via the golfing scheme
In this subsection, we construct the low-rank part of the dual certificate via the golfing scheme [25, 12]; let us start with a couple of probability bounds before introducing the scheme. The first one is a direct consequence of [75, Theorem 2.1]: If $d$th order tensors $\mathbf{X}\in\mathbb{R}^{n\times n\times\dots\times n}$ and $\mathbf{Y}\in\{0,1\}^{n\times n\times\dots\times n}$ with $\mathbf{X}$ given and $\mathbb{I}(\mathbf{Y})\sim\operatorname{Bernoulli}(q)$ for $q\geq\theta_{1}\frac{\ln n}{n}$, then
@@
By applying the scalar Bernstein inequality [66, Theorem 2.8.4], we have
@@
As a result, for a sufficiently large $\kappa_{3}>0$,
@@
holds with high probability, where the last equality holds because $\operatorname{p}_{\mathbb{L}}(\mathbf{Z}_{j-1})-\mathbf{Z}\in\mathbb{L}$ and the second inequality is due to Corollary 6.9. The proof is then completed by an overall union bound. $\Box$
@@ 6.5 Dual certificate: Sparse part via the least squares method
In order to show (34), the existence of relevant $\epsilon$-nets is required. In particular, it is stated in [66, Corollary 4.2.13] that for any $\epsilon>0$, there exists an $\epsilon$-net of $\mathbb{S}^{n}$ whose cardinality is no more than $(1+\frac{2}{\epsilon})^{n}$. $\Box$
@@
This, together with [66, Proposition 2.5.2], further implies that
@@
The result then follows directly by combining the above with [62, Lemma 1] and [62, Theorem 1]. $\Box$
@@
where $\mathbb{E}_{k}$’s are the $\epsilon$-nets used in (34). For any given $\mathbb{I}(\mathbf{S})$, the nonzero entries of $\mathbf{E}$ are i.i.d. symmetric Bernoulli random variables (taking $\pm 1$ with equal probability), according to Assumption 6.1. Hence, by Hoeffding’s inequality [66, Theorem 2.2.2], we have
@@
For any given $\mathbb{I}(\mathbf{S})$, the nonzero entries of $\mathbf{E}$ are i.i.d. symmetric Bernoulli random variables. By the two-sided version of Hoeffding’s inequality [66, Theorem 2.2.5],
@@ 6.6 Putting everything together
We arrive at the following result by combining the results of previous two subsections with an overall union bound.
@@
We are now ready to prove Theorem 6.2.
Proof of Theorem 6.2. By letting $\delta$ to be, e.g., $\frac{1}{3}$ in Corollary 6.8, we have $\|\operatorname{p}_{\mathbb{L}}\operatorname{p}_{\mathbb{I}(\mathbf{S})}\|<\frac{1}{2}$ with high probability. Besides, $\lambda=\frac{1}{\sqrt{n_{d}}}\leq\frac{1}{\sqrt{2}}<1$ since $n_{d}\geq 2$ is assumed. Moreover, Proposition 6.18 implies that there exists a $\mathbf{D}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$ satisfying (27) with high probability. By combining the above facts with Lemma 6.4, it follows that $(\mathbf{L},\mathbf{S})$ is the unique optimal solution of (24) with high probability, i.e., (24) exactly recovers $\mathbf{L}$ and $\mathbf{S}$ with high probability. $\Box$
@@ 7 Concluding remarks
In this paper, we systematically study the decomposability and the subdifferential of the tensor nuclear norm. We show that the tensor nuclear norm is decomposable over a pair of subspaces that have at least two disjoint modes, naturally generalizing the result for the matrix case. The same property holds for the tensor spectral norm as well. Based on the decomposability, we propose novel inclusions of the subdifferential of the tensor nuclear norm. They strictly enlarge the inclusion proposed in [72, Theorem 1], the only known subdifferential inclusion for tensors of an arbitrary order. While various bounds for the subdifferential of the tensor nuclear norm and several interesting examples suggest that there is no general way to explicitly characterize the subdifferential of the tensor nuclear norm in general as in the matrix case, we examine subgradients of the tensor nuclear norm in all relevant subspaces and estimate their inner and outer approximations.
@@
In fact, the conjecture holds true if the factor $\ln^{d-2}{n_{d}}$ in Corollary 6.9 can be removed. This has already been conjectured similarly in [75, Section 11]. If Conjecture 7.1 is true, then the conditions for exact recovery of the tensor robust PCA of every order, including the matrix case, would become identical.
@@
Apart from practical applications, our developments may facilitate the study of a variety of other more in-depth properties and implications of the tensor nuclear norm. For example, they may help to deduce the $C^{2}$-cone reducibility [6, Definition 3.135], subdifferential metric subregularity [21, Section 3.H], and twice epi-differentiability [59, Definition 13.6(b)] of the tensor nuclear norm; see, e.g., [18, Proposition 3.2][74, Proposition 11] and [18, Proposition 3.8], and [28, Corollary 3.6] and [48], respectively for their matrix counterparts. Besides, the developments can also be useful to understand the neural collapse [53] in training tensor-based neural networks and the relations between various nuclear-norm-regularized tensor optimization problems and their Burer-Monteiro reparameterizations [10, 11]; see, e.g., [76, Appendix C] and [39, 8, 47, 52], respectively for their matrix counterparts.
@@ Acknowledgments
This work was partially supported by the National Natural Science Foundation of China [Grants 72394364, 72394360, and 72171141]. The authors would like to thank the two anonymous reviewers for their insightful comments that have helped to improve this paper from its original version. In particular, the authors are indebted to the reviewer who has brought the references [60, 57, 27, 65, 64, 38] to their attention and suggested the last part of Lemma 3.3. The first author would like to express his sincere gratitude to Professor Simai He for his generous subsidy support when he was studying at School of Information Management and Engineering, Shanghai University of Finance and Economics, where part of the work in this paper was carried out. The first author would also like to thank his current advisor, Professor Anthony Man-Cho So, for the continued encouragement and support in pursuing this project.
@@ Appendix A Computing tensor spectral norms
Since $\mathbf{Z}+\mathbf{X}(t)$ is a subtensor of $\mathbf{U}(t)$, it is easy to see (or by [40, Theorem 3.1]) that
@@
Proof. Since $\mathbf{X}(t)$ is a symmetric tensor, it follows from Banach’s theorem (see e.g., [2] and [15, Corollary 4.2]) that
@@
We remark that $\bigl\|\mathbf{X}(\frac{1}{\sqrt{3}})\bigr\|_{\sigma}$ has been calculated in [23, Lemma 6.2] and can be easily extended to $\bigl\|\mathbf{X}(t)\bigr\|_{\sigma}$ for any $t$. $\bigl\|\mathbf{Z}+\mathbf{X}(-1)\bigr\|_{\sigma}$ has also been calculated in [23, Lemma 6.1]. In fact, the tensor $\mathbf{Z}+\mathbf{X}(-1)$ is known as an orthogonal tensor in the literature; see [41, Theorem 3.5] for more details. The result of $\bigl\|\mathbf{Z}+\mathbf{X}(t)\bigr\|_{\sigma}$ in Lemma A.3 provides a better understanding of the generalization of such a tensor.
@@ Appendix C Subdifferential of the nuclear norm for fourth-order tensors
The following result generalizes [71, Lemma 1] from $d=3$ to $d=4$.